OpenStudy (anonymous):

use the elimination method to solve the system of equations. a+7b=16 a+3b=4 possible answers. there is one solution, it is _. b. there are infinitely many solutions. C. there is no solution

OpenStudy (anonymous):

First set: a=16-7b "Just move over 7b on the upper" Then use a in the second and insert "a" into the second equation and solve b When you know b, you can insert b in the first and figure out a ;) I hope you understand, please write back if you don't :)

OpenStudy (anonymous):

so the answer is a=16-7b?

OpenStudy (anonymous):

Correct at first, now set a, into the other equation like this: (16-7b)+3b=4 And solv b :)

OpenStudy (anonymous):

so would the correct answer be 12+21b?